OpenStudy (anonymous):

Solve. -2/5 x - 9 < 9/10

OpenStudy (anonymous):

\[-\frac{ 2 }{ 5 }x-9 <\frac{ 9 }{ 10 }\] \[-\frac{ 2 }{ 5 }x <\frac{ 99 }{ 10}\] \[x >\frac{ 99 }{ 10 }\div \frac{ -2 }{ 5 }\] \[x >-\frac{ 99 }{ 4 }\]

OpenStudy (anonymous):

Thank you so much!!